OpenStudy (gamenerd123):

Which expression is equal to √5/11 ?

OpenStudy (kittiwitti1):

\[\frac{\sqrt{5}}{11}\]This? Also, what are your options?

OpenStudy (gamenerd123):

Which expression is equal to 511√511 ? 25/11 5√11/11 5/√11 √55/11

OpenStudy (gamenerd123):

Ignore the question

OpenStudy (gamenerd123):

Those are the options

OpenStudy (kittiwitti1):

\(511\sqrt{511}\)?

OpenStudy (gamenerd123):

5/√11

OpenStudy (gamenerd123):



OpenStudy (kittiwitti1):

So you want to find another form of \(\large{\frac{5}{\sqrt{11}}}\)?

OpenStudy (gamenerd123):

Yes, which expression is equal to it

OpenStudy (kittiwitti1):

Hmm, let's see. These are your options, yes? 1) \(\Large{\frac{25}{11}}\) 2) \(\Large{\frac{5\sqrt{11}}{11}}\) 3) \(\Large{\frac{5}{\sqrt{11}}}\) 4) \(\Large{\frac{\sqrt{55}}{11}}\)

OpenStudy (welshfella):

multiply top and bottom of the fraction by \[\sqrt{11}\]

OpenStudy (kittiwitti1):

I think there is more than one option as the answer if I got the options right... are you allowed to pick more than one choice? @Gamenerd123

OpenStudy (gamenerd123):

Find the Square root of 11 and then multiply? or just multiply by 11?

OpenStudy (gamenerd123):

Nope, just one

OpenStudy (kittiwitti1):

Hmm, that's weird, because the original \(\Large{\frac{5}{\sqrt{11}}}\) is also one of the options... But what @welshfella said is correct. You should do this:\[\frac{5}{\sqrt{11}}\times\frac{\sqrt{11}}{\sqrt{11}}\]

OpenStudy (welshfella):